# Probing Visual Language Priors in VLMs

### v3

@@ A.3 Comparisons to other datasets
A.3.1 Compare to Winoground (Thrush et al., 2022)
@@
A.3.2 Compare to Whoops! (Bitton-Guetta et al., 2023a)
@@
A.3.3 Compare to HallusionBench (Guan et al., 2023)
@@ B.1 RLHF for VLM
RL Fine-Tunning Phrase
@@ B.2 Image DPO and RLHF
$\mathcal{L}(\pi_{\theta};\pi_{\text{ref}})\leq-\mathbb{E}_{(Q,I_{w},I_{l},A)\sim S}\left[\log\sigma(\alpha\frac{\pi_{\theta}(A\mid Q,I_{w})}{\pi_{\theta_{\text{ref}}}(A\mid Q,I_{w})}-\alpha\frac{\pi_{\theta}(A\mid Q,I_{l})}{\pi_{\theta_{\text{ref}}}(A\mid Q,I_{l})})\right]$ (13)
@@ C.1 VLM self-guided data generation
To provide a better understanding of our generated QIAs, we randomly sampled and listed some examples of the generated QIA data, as shown in Figures 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, and 33.
@@ C.2 Image DPO data preparation and training details
This section details the construction of good-bad question-image-answer (QIA) pairs ($I_{w}$, $I_{l}$) based on the QIAs generated by the pipeline described in Appendix C.1. In brief, the data generation pipeline outlined in Appendix C.1 utilizes VLMs in conjunction with pre-trained image models to generate or modify images and create corresponding question-answer pairs. This process results in a collection of QIA triplets, as illustrated in the Figures 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, and 33.

### v4

@@ A.4 Comparisons to other datasets
A.4.1 Compare to Winoground (Thrush et al., 2022)
@@
A.4.2 Compare to Whoops! (Bitton-Guetta et al., 2023a)
@@
A.4.3 Compare to HallusionBench (Guan et al., 2023)
@@ A.5 Human Study
Figure 14: Failure cases analysis. We list six typical failure cases generated by VLMs on ViLP.
A.6 More failure cases analysis
This subsection provides several typical failure cases observed in the inferred results of VLMs on ViLP. We identified several consistent failure patterns, illustrated in Figure 14. For deeper analysis, we also prompted the VLMs to output their reasoning before finalizing answers, allowing us to better understand how these errors arise. Below, we summarize the failure patterns in the same order as they appear in the figure, from left to right and top to bottom.
- Shape recognition can fail in VLMs, causing them to revert to priors rather than accurately interpreting visual input.
- Models sometimes struggle to count accurately. Instead of performing an actual count, they default to relying on learned priors to estimate quantities.
- Models may refuse to accept visual information that contradicts their learned priors, whereas humans can comprehend hypothetical scenarios. For instance, the model recognizes the city as Beijing but rejects the correct answer because it expects the Eiffel Tower to be in Paris.
- Sometimes the model overly relies on memorized proverbs, resulting in predictions that align with these proverbs rather than the actual content of the input image.
- For images with creative concepts, the model may overly rely on its learned priors. As illustrated, a common prior is that microscopes are used to view organisms, leading the model to answer “microscope” rather than identifying the creatively depicted stapler.
- For images with blended features, the model may rely mostly on text input while overlooking the visual cues. As illustrated, the VLM heavily depends on textual input leading to saxophone as the answer.
@@ B.1 RLHF for VLM
RL Fine-Tuning Phrase
@@ B.2 Image DPO and RLHF
$\mathcal{L}(\pi_{\theta};\pi_{\text{ref}})\leq-\mathbb{E}_{(Q,I_{w},I_{l},A)\sim S}\left[\log\sigma(\alpha\frac{\pi_{\theta}(A\mid Q,I_{w})}{\pi_{{\text{ref}}}(A\mid Q,I_{w})}-\alpha\frac{\pi_{\theta}(A\mid Q,I_{l})}{\pi_{{\text{ref}}}(A\mid Q,I_{l})})\right]$ (13)
@@ C.1 VLM self-guided data generation
To provide a better understanding of our generated QIAs, we randomly sampled and listed some examples of the generated QIA data, as shown in Figures 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, and 35.
@@ C.2 Image DPO data preparation and training details
This section details the construction of good-bad question-image-answer (QIA) pairs ($I_{w}$, $I_{l}$) based on the QIAs generated by the pipeline described in Appendix C.1. In brief, the data generation pipeline outlined in Appendix C.1 utilizes VLMs in conjunction with pre-trained image models to generate or modify images and create corresponding question-answer pairs. This process results in a collection of QIA triplets, as illustrated in the Figures 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, and 35.